solving a polynomial equation.
4y^3-7y^2-16y+28=0

Answer by edjones(8007)

4y^3-7y^2-16y+28=0
y^2(4y-7)-4(4y-7)=0
(y^2-4)(4y-7)=0
(y+2)(y-2)(4y-7)=0
y=-2, y=2, y=7/4
.
Ed